A man is 30 years older than the wife and there childs age is average of the mother's age if the total age is 105 find the following
the age of the father
the age of the mother
the age of the child

Answer by Timnewman(323)

Hi,
Let the mother's age be "x"
Then,
Father's age =x+30
Also,
Child's age =x/2
But the sum of their age is 105
Then,
x+x+30+(x/2)=105
2x+(x/2)=105-30
2x+(x/2)=75.
multiply althrough by 2
4x+x=150
5x=150
x=150/5
x=30
The mother's age is 30yrs
Father's age=x+30
=30+30
=60yrs
Also,Child's age =x/2
=30/2
=15yrs
The mother's age =30yrs
Father's age=60yrs
Child's age=15yrs.
HOPE THIS HELPS,....
Best regards
---TIMNEWMAN.